OpenStudy (anonymous):

\[Sketch~the~graph~of~~~y=xe^{-x^2} ~~~for~~~~x>0.\]\[ ~Find~the~area~contained~\between~this~graph~and~the~positive~x-axis\]

OpenStudy (anonymous):

Hey, tiaph, I'm going. I'll see the post when you write it! :)

OpenStudy (anonymous):

I tried sketching the graph out for u but i couldnt, so here is a link, http://www.wolframalpha.com/input/?i=x*e^%28-x^2%29 since the question asks for the graph for x>0 take note to draw only that part of the graph u see in the link. To find the area, u neecd to integrate the function. \[\int\limits_{0}^{\infty}x \times e ^{-x ^{2}} \] Integrating it, u get. \[-\frac{1}{2 e ^{x ^{2}}}+c \], where c is a arbitrary constant. Proceed to solve the definite integrals. \[- \frac{1}{2e ^{\infty ^{2}}} -( - \frac{1}{2e ^{0^{2}}}) = 0 + \frac{1}{2} =\frac{1}{2}\] e^ infinity is infinity. 1divided by infinity can be approximated to be 0.

OpenStudy (anonymous):

Hm... Ok thanks!!! :)

OpenStudy (anonymous):

Np!